OpenStudy (anonymous):

There is no solution to the system below. x + 3y = 7 2x + 6y = 9 True False

OpenStudy (anonymous):

sove the system by your favorite method :)

OpenStudy (anonymous):

and then you'll see if there is no solution

OpenStudy (anonymous):

I got TRUEEEEEEEEEEEEEEEE

OpenStudy (anonymous):

I dont think so :(

OpenStudy (anonymous):

oops:(((

OpenStudy (anonymous):

here is another method since this is a simple 2 equations and 2 variables x + 3y = 7 2x + 6y = 9 solve each equation for y y = -(1/3)*x+7/3 y = -(1/3)*x+3/2 as you can see it will be in the y=mx + b form the m aka slopes are the same. which means they dont meet ever. and the bs are different. therefore no solution. When Ms are same and Bs are different.

OpenStudy (anonymous):

If you're familiar with determinants, you can calculate \[\begin{vmatrix} 1 & 3 \\ 2 & 6\end{vmatrix} = (1)(6)-(3)(2) = 0,\] so there is no solution.